OpenStudy (anonymous):

find the vertex f(x)= x^2-2x-8

OpenStudy (anonymous):

first we derive f(x)

OpenStudy (anonymous):

f´(x) = 2x - 2

OpenStudy (anonymous):

then 2x -2 = 0 x = 1

OpenStudy (anonymous):

we evaluate x = 1 at original equation

OpenStudy (anonymous):

f(x) = y = 1^2 - 2 (1) - 8 = 1 - 2 - 8 = -9

OpenStudy (anonymous):

vertex i located at ( 1 ; -9)

OpenStudy (anonymous):

It was a pleaure to help you